# The evolution of the galaxy stellar mass function and star formation rates in the COLIBRE simulations from redshift 17 to 0

### v1

@@ 1 Introduction
At low redshifts ($z\lesssim 1$), the GSMF has been tightly constrained by large spectroscopic and photometric surveys, such as the Galaxy And Mass Assembly Survey (GAMA, Driver et al., 2011; Baldry et al., 2012; Liske et al., 2015; Driver et al., 2022), the Cosmic Evolution Survey (COSMOS, Scoville et al., 2007; Davidzon et al., 2017; Weaver et al., 2022), and the Deep Extragalactic VIsible Legacy Survey (DEVILS, Davies et al., 2018; Thorne et al., 2021). These surveys have constructed galaxy samples complete down to stellar masses of $M_{*}\sim 10^{7}-10^{8}~\mathrm{M_{\odot}}$, encompassing up to $\sim 10^{6}$ objects over areas of sky of $\sim 1$ to $10^{2}$ deg2. Multi-wavelength photometry – spanning (far-)UV to (far-)IR wavelengths – is typically available in $\approx 10-30$ bands per galaxy. These extensive data enable robust stellar mass determination via spectral energy distribution (SED) fitting, with systematic uncertainties generally within $\approx 0.15$ dex (e.g. Robotham et al., 2020). Furthermore, very recently, Xu et al. (2025) probed the faint end of the $z\approx 0$ GSMF down to $\sim 10^{6}~\mathrm{M_{\odot}}$ by applying the statistical Photometric Objects Around Cosmic-webs (PAC) method to the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al. 2016) survey data. PAC combines a large-scale (but relatively shallow) spectroscopic sample with deep photometric data to statistically associate faint galaxies (lacking spectroscopy) with nearby spectroscopic tracers, enabling redshift determination and stellar mass estimates for galaxies too faint to be targeted spectroscopically.
@@
L050m5 $50$ $2.3\times 10^{5}$ $3.0\times 10^{5}$ $1504^{3}$ $4\times 1504^{3}$ $\mathrm{min(0.35\,pkpc,0.9\,ckpc)}$ thermal $2.5$
@@
The unprecedented IR sensitivity and resolution of JWST have prompted a re-evaluation of how stellar mass accumulated in the high-redshift Universe. In particular, JWST observations at $z\gtrsim 4$ reveal a more rapid growth of massive galaxies than previously anticipated (e.g. Bunker et al., 2023; Weibel et al., 2024; Wang et al., 2024; Carnall et al., 2024), suggesting that galaxy formation efficiencies $\gtrsim 30$ per cent may be necessary to reach the observed stellar masses ($M_{*}\gtrsim 10^{10}~\mathrm{M_{\odot}}$) within the age of the Universe at the observed redshift (e.g. Harikane et al., 2023; Casey et al., 2024). The SEDs of these massive galaxies often indicate significant dust content and suggest that some of these galaxies are already quenched (i.e. no longer actively forming stars) at those high redshifts – a challenge that most theoretical models struggle to reproduce (e.g. Valentino et al., 2023; Baker et al., 2025b; Weibel et al., 2025b; de Graaff et al., 2025; Lagos et al., 2025). Some of these conclusions, however, may be affected by significant uncertainties in stellar mass estimates (within $\approx 0.5$ dex, e.g. Cochrane et al. 2025, Wang et al. 2025), cosmic variance due to the limited surface areas probed by high-$z$ surveys, and the reliance on photometric redshifts for most galaxies.
@@
In this work, we use the new colibre cosmological hydrodynamical simulations of galaxy formation (Schaye et al., 2025; Chaikin et al., 2025; Huško et al., 2025a) to investigate the buildup of stellar mass in the Universe across cosmic time. We compare simulation predictions with observational data from $z=0$ up to the highest redshift with available observations ($z\approx 12$ for the GSMF) and make predictions out to $z=17$. Unlike previous galaxy formation models applied to cosmological volumes of comparable size, colibre includes a multiphase interstellar medium (ISM) with non-equilibrium chemistry for hydrogen and helium (Ploeckinger et al., 2025), employs a detailed model for the formation and evolution of dust grains coupled to the chemistry (Trayford et al., 2025), and suppresses spurious transfer of energy from dynamically-hot dark matter (DM) to dynamically-cooler stars by using four times as many DM particles as baryonic particles. In addition to the GSMF, we examine several other statistics closely tied to stellar mass evolution: the cosmic star formation rate and stellar mass densities, the stellar-to-halo mass relation (SHMR), the star-forming main sequence (SFMS), the galaxy quenched fraction – stellar mass relation, and the number density of massive quiescent galaxies in the high-redshift Universe. The paper is structured as follows: Section 2 details the simulation methodology, Section 3 presents our findings, and Section 4 summarizes our key conclusions.
@@ 2.1 The COLIBRE simulations
In this work, we use the colibre simulations at m7 resolution with comoving volumes of $400^{3}$ and $200^{3}$ cMpc3, at m6 resolution with a volume of $200^{3}$ cMpc3, and at m5 resolution [^1] At the moment of writing, the m5 simulation in the $50^{3}$ cMpc3 volume is at $z=2.5$. Therefore, unless stated otherwise, results at m5 resolution are shown from the smaller $25^{3}$ cMpc3 volume for $z<2.5$ and from the $50^{3}$ cMpc3 volume for $z\geq 2.5$. with $50^{3}$ and $25^{3}$ cMpc3. Additionally, to assess the separate convergence with resolution and with cosmological volume (see Appendix A), we use smaller volumes at m7 and m6 resolutions: $100^{3}~\mathrm{cMpc}^{3}$, $50^{3}~\mathrm{cMpc}^{3}$, and $25^{3}~\mathrm{cMpc}^{3}$. Table 1 provides a summary of all simulations used in this study. Unless specified otherwise, we use the largest available simulation at each resolution.
@@ 2.1.1 The subgrid model
Following Nobels et al. (2024), star-forming gas particles are identified using the gravitational instability criterion. To assign an SFR to these particles, we apply the Schmidt (1959) law, assuming a fixed star formation efficiency per free-fall time of $\varepsilon=0.01$. Newly formed stellar particles represent simple stellar populations characterized by a Chabrier (2003) IMF. Stellar particles enrich their surrounding gas with metals through six chemical enrichment channels: AGN stars, type-Ia SNe, type-II SNe, neutron star mergers, common envelope jet SNe, and collapsars (see Correa et al., in preparation, for further details).
@@
colibre also includes three early stellar feedback processes: H ii regions, stellar winds, and radiation pressure (Benítez-Llambay et al., in preparation). These processes are implemented stochastically, similarly to the energy feedback from SNe.
@@
Supermassive black holes (SMBHs) are represented by black hole (BH) particles, seeded at high redshift by converting the densest gas particle in a halo into a BH particle. Seeding occurs when the halo mass, calculated on-the-fly using the Friends-of-Friends (FoF) algorithm, exceeds a threshold of $5\times 10^{10}~\mathrm{M_{\odot}}$ at m7 resolution or $10^{10}~\mathrm{M_{\odot}}$ at m6 and m5 resolutions, provided the FoF halo does not already contain a BH particle. The mass accretion rate onto BHs is computed using the Bondi-Hoyle-Lyttleton formula, with turbulence and vorticity corrections from Krumholz et al. (2006). Dynamical friction and BH-BH mergers are modelled following the prescriptions from Bahé et al. (2022) and Schaye et al. (2025).
@@
At each resolution, the colibre simulations include models with purely thermal AGN feedback and hybrid AGN feedback (combining kinetic jets and thermal energy injections), with the largest colibre volumes available only for the thermal models (see table 2 in Schaye et al., 2025). The primary focus of this work is on the colibre simulations with purely thermal AGN feedback, though we provide a comparison between the hybrid and thermal models in $\S$3.8.
@@
Thermal AGN feedback is implemented deterministically, following Booth & Schaye (2009), but with BH particles heating their neighbouring gas by a temperature increment proportional to the BH (subgrid) mass, $\Delta T_{\rm AGN}\propto m_{\rm BH}$, rather than by a fixed temperature increment. The BH mass-dependent heating ensures that the frequency of AGN feedback events remains independent of BH mass at a fixed Eddington fraction, and that at the highest colibre resolution, m5, the BH mass does not increase too steeply with galaxy stellar mass (see Chaikin et al. 2025 for details).
@@
The hybrid AGN model detailed in Huško et al. (2025a), additionally tracks the evolution of BH spin and distinguishes between three accretion states: the thick, thin and slim discs. The spin of each BH particle evolves over time due to gas accretion, BH-BH mergers, and jet-induced spindown. The kinetic jet feedback mode is implemented deterministically: once a BH particle has accreted a mass whose energy equivalent exceeds that of a single jet event, two neighbouring gas particles are kicked in opposite directions, nearly aligned with the BH spin (within $7.5$ degrees of the spin axis), with the jet kick velocity proportional to the square root of the BH (subgrid) mass, such that the injected energy scales linearly with BH mass. The thermal channel for energy injection is implemented identically to that in the purely thermal AGN feedback model, but with a spin-dependent radiative efficiency for the thin disc, while separate terms are used to represent the effects of winds in the thick and slim disc states. The amounts of energy available for injection through the thermal and kinetic jet channels are determined based on the BH’s instantaneous accretion rate, expressed as a fraction of the Eddington rate, and its spin. In both the thermal and hybrid AGN feedback models, the BH accretion rate is capped at 100 times the Eddington rate, allowing for super-Eddington accretion events.
@@ 2.1.2 Calibration
For the colibre models incorporating hybrid AGN feedback, AGN feedback parameters were manually tuned at each resolution to match the same calibration data as for the thermal models, as well as the $z=0.2$ AGN bolometric luminosity function (see Huško et al., 2025a). Among the SN feedback parameters, only the normalization of the $\Delta T_{\rm SN}(\rho_{\rm gas})\propto\rho_{\rm gas}^{2/3}$ scaling was adjusted, while all other SN feedback parameters retained the same values as in the colibre models with thermal AGN feedback at the corresponding resolution.
@@ 2.2 Halo finding and calculation of galaxy properties
We identify subhaloes using HBT-HERONS [^3] https://hbt-herons.strw.leidenuniv.nl/ (Forouhar Moreno et al., 2025), which is an updated version of the Hierarchical Bound Tracing algorithm (HBT) originally presented by Han et al. (2018). HBT-HERONS identifies structures by applying an iterative unbinding algorithm to particles within spatial FoF groups, starting at the highest-redshift simulation output. Once self-bound objects are identified, their particles are tracked across subsequent simulation outputs to trace the evolution of these objects forward in time. This tracking improves the identification of satellite subhaloes, as particles remain associated with their subhaloes even after the subhalo is accreted into a larger host halo. Each candidate substructure undergoes additional checks for self-boundness and phase-space consistency to determine its status – whether it remains resolved, has merged with another substructure, or has been disrupted.
@@
  • We use the stellar mass computed by summing up the masses of individual stellar particles that are gravitationally bound to the subhalo and within 3D spherical apertures of radius 50 pkpc. The aperture is centred on the position of the particle with the largest binding energy within the subhalo. We chose the 50 pkpc aperture because de Graaff et al. (2022) found it best reproduces masses inferred from virtual observations of the eagle simulations. We discuss the impact of the aperture choice on the GSMF in Appendix B, finding a noticeable effect only for stellar masses $M_{*}\gtrsim 10^{11}~\mathrm{M_{\odot}}$.
  • The SFR is calculated by summing the instantaneous SFRs of gas particles that are both bound to the subhalo and within a 50 pkpc 3D spherical aperture. Additionally, in Appendix C, we explore the impact of using 10- and 100-Myr averaged SFRs instead of instantaneous SFRs, finding generally minor differences at all redshifts.
@@
We use SOAP outputs at integer and half-integer redshifts from $z=20$ to $z=0$, and additionally at $z=0.1$.
@@ 2.3 Assumptions in the analysis
Figure 1: Evolution of the galaxy stellar mass function (GSMF) from $z=0$ (top left) to $z=4$ (bottom right) in the colibre m7, m6, and m5 simulations using the largest cosmological volumes available (see Table 1), with solid dark-red, orange, and light-blue curves representing each resolution, respectively. To highlight resolution and volume limitations, the solid lines become dotted at stellar masses where galaxies are poorly resolved ($M_{*}<10$ times the mean initial baryonic particle mass) or where the number of galaxies per $0.2$ dex mass bin is less than $10$. A lognormal scatter with a standard deviation given by equation (1) is added to galaxy stellar masses to account for Eddington bias. For reference, the (full extent of the) horizontal black error bar in each panel indicates the typical systematic uncertainty in stellar mass measurements, as given by equation (2). For comparison, we include a compilation of observational data from Leja et al. (2020), Stefanon et al. (2021), Thorne et al. (2021), Driver et al. (2022), Weaver et al. (2023b), Weibel et al. (2024), Navarro-Carrera et al. (2024), and Shuntov et al. (2025) (black symbols). colibre is consistent with the selected observational data, given the systematic uncertainties on observationally inferred stellar masses, and achieves excellent convergence with resolution at all explored redshifts.
@@
Throughout this work, we assume a Chabrier (2003) IMF. All observational data based on Salpeter (1955) or Kroupa (2001) IMFs that we include for comparison with the simulations are converted to the Chabrier (2003) IMF using the conversion factors for SFRs and stellar masses provided by Madau & Dickinson (2014). Additionally, we account for differences in the assumed value of the dimensionless Hubble parameter $h$ between the cosmologies used in the comparison data and in colibre.
@@
Random errors in observationally inferred stellar masses, arising from the application of SED-fitting algorithms to broad-band photometric data and uncertainties in SPS modelling, are estimated to be $\approx 0.1-0.2$ dex at $z\lesssim 1$ (e.g. Gallazzi & Bell, 2009; Taylor et al., 2011; Speagle et al., 2014; Mobasher et al., 2015; Robotham et al., 2020; Jones et al., 2022; Lee et al., 2025), and $\approx 0.2-0.3$ dex or higher at $z\gtrsim 1$ (e.g. Conroy et al., 2009; Santini et al., 2023; Pacifici et al., 2023; Cochrane et al., 2025). These random errors affect the shape of the observed GSMF through the Eddington (1913) bias. Due to the steep slope of the GSMF at $M_{*}\gtrsim 10^{11}~\mathrm{M_{\odot}}$, there are significantly more low-mass galaxies than high-mass ones, leading to a greater upward scatter of random errors in stellar mass of low-mass objects compared to the downward scatter of high-mass objects. As we will explore in Appendix D, this effect has an impact on any statistic related to galaxy stellar mass, including the GSMF, the SFMS, the galaxy quenched fraction, and the number density of massive quiescent galaxies.
@@
Figure 2: As Fig. 1, but for higher redshifts, ranging from $z=5$ (top left) to $z=17$ (bottom right). Black symbols represent observational data from Stefanon et al. (2021), Weaver et al. (2023b), Weibel et al. (2024), Navarro-Carrera et al. (2024), Shuntov et al. (2025), and Harvey et al. (2025). colibre again shows very good convergence with resolution and is consistent with the observational data.

### v2

@@ 1 Introduction
At low redshifts ($z\lesssim 1$), the GSMF has been tightly constrained by large spectroscopic and photometric surveys, such as the Galaxy And Mass Assembly Survey (GAMA, Driver et al., 2011; Baldry et al., 2012; Liske et al., 2015; Driver et al., 2022), the Cosmic Evolution Survey (COSMOS, Scoville et al., 2007; Davidzon et al., 2017; Weaver et al., 2022), the Deep Extragalactic VIsible Legacy Survey (DEVILS, Davies et al., 2018; Thorne et al., 2021), and, very recently, the Cosmic Dawn Survey (Euclid Collaboration et al., 2025b, c). These surveys, some of which also probe higher redshifts, have constructed galaxy samples at $z\lesssim 1$ that are complete down to stellar masses of $M_{*}\sim 10^{7}-10^{9}~\mathrm{M_{\odot}}$ (depending on redshift), encompassing up to $\sim 10^{6}$ objects over areas of sky of $\sim 1$ to $10^{2}$ deg2. Multi-wavelength photometry – spanning (far-)UV to (far-)IR wavelengths – is typically available in $\approx 10-30$ bands per galaxy. These extensive data enable robust stellar mass determination via spectral energy distribution (SED) fitting, with systematic uncertainties generally within $\approx 0.15$ dex at $z\approx 0$ (e.g. Robotham et al., 2020). Furthermore, very recently, Xu et al. (2025) probed the faint end of the $z\approx 0$ GSMF down to $\sim 10^{6}~\mathrm{M_{\odot}}$ by applying the statistical Photometric Objects Around Cosmic-webs (PAC) method to the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al. 2016) survey data. PAC combines a large-scale (but relatively shallow) spectroscopic sample with deep photometric data to statistically associate faint galaxies (lacking spectroscopy) with nearby spectroscopic tracers, enabling redshift determination and stellar mass estimates for galaxies too faint to be targeted spectroscopically.
@@
L100m5 $100$ $2.3\times 10^{5}$ $3.0\times 10^{5}$ $3008^{3}$ $4\times 3008^{3}$ $\mathrm{min(0.35\,pkpc,0.9\,ckpc)}$ thermal $4.6$
L050m5 $50$ $2.3\times 10^{5}$ $3.0\times 10^{5}$ $1504^{3}$ $4\times 1504^{3}$ $\mathrm{min(0.35\,pkpc,0.9\,ckpc)}$ thermal $0.9$
@@
The unprecedented IR sensitivity and resolution of JWST have prompted a re-evaluation of how stellar mass accumulated in the high-redshift Universe. In particular, JWST observations at $z\gtrsim 4$ reveal a more rapid growth of massive galaxies than previously anticipated (e.g. Bunker et al., 2023; Weibel et al., 2024; Carnall et al., 2024; Wang et al., 2025), suggesting that galaxy formation efficiencies $\gtrsim 30$ per cent may be necessary to reach the observed stellar masses ($M_{*}\gtrsim 10^{10}~\mathrm{M_{\odot}}$) within the age of the Universe at the observed redshift (e.g. Harikane et al., 2023; Casey et al., 2024). The SEDs of these massive galaxies often indicate significant dust content and suggest that some of these galaxies are already quenched (i.e. no longer actively forming stars) at those high redshifts – a challenge that most theoretical models struggle to reproduce (e.g. Valentino et al., 2023; Baker et al., 2025a; Weibel et al., 2025b; de Graaff et al., 2025; Lagos et al., 2025). Some of these conclusions, however, may be affected by significant uncertainties in stellar mass estimates (within $\approx 0.5$ dex, e.g. Cochrane et al. 2025, Wang et al. 2025), cosmic variance due to the limited surface areas probed by high-$z$ surveys, and the reliance on photometric redshifts for most galaxies.
@@
In this work, we use the new colibre cosmological hydrodynamical simulations of galaxy formation (Schaye et al., 2026; Chaikin et al., 2026; Huško et al., 2026) to investigate the buildup of stellar mass in the Universe across cosmic time. We compare simulation predictions with observational data from $z=0$ up to the highest redshift with available observations ($z\approx 12$ for the GSMF) and make predictions out to $z=17$. Unlike previous galaxy formation models applied to cosmological volumes of comparable size, colibre includes a multiphase interstellar medium (ISM) with non-equilibrium chemistry for hydrogen and helium (Ploeckinger et al., 2025), employs a detailed model for the formation and evolution of dust grains coupled to the chemistry (Trayford et al., 2026), and suppresses spurious transfer of energy from dynamically-hot dark matter (DM) to dynamically-cooler stars by using four times as many DM particles as baryonic particles. In addition to the GSMF, we examine several other statistics closely tied to stellar mass evolution: the cosmic star formation rate and stellar mass densities, the stellar-to-halo mass relation (SHMR), the star-forming main sequence (SFMS), the galaxy quenched fraction – stellar mass relation, and the number density of massive quiescent galaxies in the high-redshift Universe. The paper is structured as follows: Section 2 details the simulation methodology, Section 3 presents our findings, and Section 4 summarizes our key conclusions.
@@ 2.1 The COLIBRE simulations
In this work, we use the colibre simulations at m7 resolution with comoving volumes of $400^{3}$ and $200^{3}$ cMpc3, at m6 resolution with a volume of $200^{3}$ cMpc3, and at m5 resolution with $100^{3}$, $50^{3}$, and $25^{3}$ cMpc3. Additionally, to assess the separate convergence with resolution and with cosmological volume (see Appendix A), we use smaller volumes at m7 and m6 resolutions: $100^{3}~\mathrm{cMpc}^{3}$, $50^{3}~\mathrm{cMpc}^{3}$, and $25^{3}~\mathrm{cMpc}^{3}$. Table 1 provides a summary of all simulations used in this study. Unless specified otherwise, we use the largest available simulation at each resolution [^1] At the time of writing, the m5 simulation in the $100^{3}$ ($50^{3}$) cMpc3 volume is at $z=4.6$ ($z=0.9$). Therefore, unless stated otherwise, results at m5 resolution are shown from the largest available volume at a given redshift, while time-evolving quantities are interpolated around $z=4.6$ and $z=0.9$, switching from L100m5 to L050m5 and from L050m5 to L025m5, respectively..
@@ 2.1.1 The subgrid model
Following Nobels et al. (2024), star-forming gas particles are identified using the gravitational instability criterion. To assign an SFR to these particles, we apply the Schmidt (1959) law, assuming a fixed star formation efficiency per free-fall time of $\varepsilon=0.01$. Newly formed stellar particles represent simple stellar populations characterized by a Chabrier (2003) IMF. Stellar particles enrich their surrounding gas with metals through six chemical enrichment channels: AGB stars, type-Ia SNe, type-II SNe, neutron star mergers, common envelope jet SNe, and collapsars (see Correa et al., 2026, for further details).
@@
colibre also includes three early stellar feedback processes: H ii regions, stellar winds, and radiation pressure (Benítez-Llambay et al., 2026). These processes are implemented stochastically, similarly to the energy feedback from SNe.
@@
Supermassive black holes (SMBHs) are represented by black hole (BH) particles, seeded at high redshift by converting the densest gas particle in a halo into a BH particle. Seeding occurs when the halo mass, calculated on-the-fly using the Friends-of-Friends (FoF) algorithm, exceeds a threshold of $5\times 10^{10}~\mathrm{M_{\odot}}$ at m7 resolution or $10^{10}~\mathrm{M_{\odot}}$ at m6 and m5 resolutions, provided the FoF halo does not already contain a BH particle. The mass accretion rate onto BHs is computed using the modified Bondi-Hoyle-Lyttleton formula, with turbulence and vorticity corrections from Krumholz et al. (2006). Dynamical friction and BH-BH mergers are modelled following the prescriptions from Bahé et al. (2022) and Schaye et al. (2026).
@@
At each resolution, the colibre simulations include models with purely thermal AGN feedback and hybrid AGN feedback (combining kinetic jets and thermal energy injections), with the largest colibre volumes available only for the thermal models (see table 2 in Schaye et al., 2026). The primary focus of this work is on the colibre simulations with purely thermal AGN feedback, though we provide a comparison between the hybrid and thermal models in $\S$3.8.
@@
Thermal AGN feedback is implemented deterministically, following Booth & Schaye (2009), but with BH particles heating their neighbouring gas by a temperature increment proportional to the BH (subgrid) mass, $\Delta T_{\rm AGN}\propto m_{\rm BH}$, rather than by a fixed temperature increment. The BH mass-dependent heating ensures that the frequency of AGN feedback events remains independent of BH mass at a fixed Eddington fraction, and that at the highest colibre resolution, m5, the BH mass does not increase too steeply with galaxy stellar mass (see Chaikin et al. 2026 for details).
@@
The hybrid AGN model detailed in Huško et al. (2026), additionally tracks the evolution of BH spin and distinguishes between three accretion states: the thick, thin and slim discs. The spin of each BH particle evolves over time due to gas accretion, BH-BH mergers, and jet-induced spindown. The kinetic jet feedback mode is implemented deterministically: once a BH particle has accreted a mass whose energy equivalent exceeds that of a single jet event, two neighbouring gas particles are kicked in opposite directions, nearly aligned with the BH spin (within $7.5$ degrees of the spin axis), with the jet kick velocity proportional to the square root of the BH (subgrid) mass, such that the injected energy scales linearly with BH mass. The thermal channel for energy injection is implemented identically to that in the purely thermal AGN feedback model, but with a spin-dependent radiative efficiency for the thin disc, while separate terms are used to represent the effects of winds in the thick and slim disc states. The amounts of energy available for injection through the thermal and kinetic jet channels are determined based on the BH’s instantaneous accretion rate, expressed as a fraction of the Eddington rate, and its spin. In both the thermal and hybrid AGN feedback models, the BH accretion rate is capped at 100 times the Eddington rate, allowing for super-Eddington accretion events.
@@ 2.1.2 Calibration
For the colibre models incorporating hybrid AGN feedback, AGN feedback parameters were manually tuned at each resolution to match the same calibration data as for the thermal models, as well as the $z=0.2$ AGN bolometric luminosity function (see Huško et al., 2026). Among the SN feedback parameters, only the normalization of the $\Delta T_{\rm SN}(\rho_{\rm gas})\propto\rho_{\rm gas}^{2/3}$ scaling was adjusted, while all other SN feedback parameters retained the same values as in the colibre models with thermal AGN feedback at the corresponding resolution.
@@ 2.2 Halo finding and calculation of galaxy properties
We identify subhaloes using HBT-HERONS [^3] https://hbt-herons.strw.leidenuniv.nl/ (Forouhar Moreno et al., 2025), which is an updated version of the Hierarchical Bound Tracing algorithm (HBT+) originally presented by Han et al. (2018). HBT-HERONS identifies structures by applying an iterative unbinding algorithm to particles within spatial FoF groups, starting at the highest-redshift simulation output. Once self-bound objects are identified, their particles are tracked across subsequent simulation outputs to trace the evolution of these objects forward in time. This tracking improves the identification of satellite subhaloes, as particles remain associated with their subhaloes even after the subhalo is accreted into a larger host halo. Each candidate substructure undergoes additional checks for self-boundness and phase-space consistency to determine its status – whether it remains resolved, has merged with another substructure, or has been disrupted.
@@
  • We use the stellar mass computed by summing up the masses of individual stellar particles that are both gravitationally bound to the subhalo and located within 3D spherical apertures of radius 50 pkpc. The aperture is centred on the position of the particle with the largest binding energy within the subhalo. We chose the 50 pkpc aperture because de Graaff et al. (2022) found it best reproduces masses inferred from virtual observations of the eagle simulations. We discuss the impact of the aperture choice on the GSMF in Appendix B, finding a noticeable effect only for stellar masses $M_{*}\gtrsim 10^{11}~\mathrm{M_{\odot}}$.
  • The SFR is calculated by summing the instantaneous SFRs of gas particles that are both bound to the subhalo and within a 50 pkpc 3D spherical aperture [^4] We verified that the results of this work are largely insensitive to whether galaxy SFRs are computed using all star-forming gas particles within a 50 kpc aperture or only the gravitationally bound particles within that same distance.. Additionally, in Appendix C, we explore the impact of using 10- and 100-Myr averaged SFRs instead of instantaneous SFRs, finding generally minor differences at all redshifts.
@@
We use SOAP outputs at integer redshifts from $z=20$ to $z=12$; at integer and half-integer redshifts from $z=12$ to $z=10$; at integer, half-, and quarter-integer redshifts from $z=10$ to $z=1$; and at lower redshifts with a step of $\Delta z=0.05$, except at $z=0.15$, where we instead use $z=0.14$ due to the structure of the colibre output redshifts.
@@ 2.3 Assumptions in the analysis
Figure 1: Evolution of the galaxy stellar mass function (GSMF) from $z=0$ (top left) to $z=4$ (bottom right) in the colibre m7, m6, and m5 simulations using the largest cosmological volumes available (see Table 1), with solid dark-red, orange, and light-blue curves representing each resolution, respectively. To highlight resolution and volume limitations, the solid lines become dotted at stellar masses where galaxies are poorly resolved ($M_{*}<10$ times the mean initial baryonic particle mass) or where the number of galaxies per $0.2$ dex mass bin is less than $10$. A lognormal scatter with a standard deviation given by equation (1) is added to galaxy stellar masses to account for Eddington bias. For reference, the (full extent of the) horizontal black error bar in each panel indicates the typical systematic uncertainty in stellar mass measurements, as given by equation (2). For comparison, we include a compilation of observational data from Leja et al. (2020), Stefanon et al. (2021), Thorne et al. (2021), Driver et al. (2022), Weaver et al. (2023b), Weibel et al. (2024), Navarro-Carrera et al. (2024), Shuntov et al. (2025), and Euclid Collaboration et al. (2025b) (black symbols). colibre is consistent with the selected observational data, given the systematic uncertainties on observationally inferred stellar masses, and achieves very good convergence with resolution at all explored redshifts.
@@
Throughout this work, we assume a Chabrier (2003) IMF. All observational data based on Salpeter (1955) or Kroupa (2001) IMFs that we include for comparison with the simulations are converted to the Chabrier (2003) IMF using the conversion factors for SFRs and stellar masses provided by Madau & Dickinson (2014). Additionally, we account for differences in the assumed value of the dimensionless Hubble parameter $h$ between cosmologies used in the comparison data, $h_{\rm obs}$, and in colibre, $h=0.681$. This is done by applying the scaling factor $(h_{\rm obs}/0.681)^{\alpha}$ to the reported observational volumes, stellar masses, and SFRs, where the exponent $\alpha$ is chosen according to each quantity’s dependence on cosmological distances.
@@
Random errors in observationally inferred stellar masses, arising from the application of SED-fitting algorithms to broad-band photometric data and uncertainties in SPS modelling, are estimated to be $\approx 0.1-0.2$ dex at $z\lesssim 1$ (e.g. Gallazzi & Bell, 2009; Taylor et al., 2011; Speagle et al., 2014; Mobasher et al., 2015; Robotham et al., 2020; Jones et al., 2022; Lee et al., 2025), and $\approx 0.2-0.3$ dex or higher at $z\gtrsim 1$ (e.g. Conroy et al., 2009; Santini et al., 2023; Pacifici et al., 2023; Cochrane et al., 2025). These random errors affect the shape of the observed GSMF through the Eddington (1913) bias, particularly at the high-mass end. Due to the steep slope of the GSMF at high $M_{*}$, there are significantly more low-mass galaxies than high-mass ones, leading to a greater upward scatter of random errors in stellar mass of low-mass objects compared to the downward scatter of high-mass objects. As we will explore in Appendix D, this effect has an impact on any statistic related to galaxy stellar mass, including the GSMF, the SFMS, the galaxy quenched fraction, and the number density of massive quiescent galaxies.
@@
Figure 2: As Fig. 1, but for higher redshifts, ranging from $z=5$ (top left) to $z=17$ (bottom right). Black symbols represent observational data from Stefanon et al. (2021), Weaver et al. (2023b), Weibel et al. (2024), Navarro-Carrera et al. (2024), Shuntov et al. (2025), Harvey et al. (2025), and Euclid Collaboration et al. (2025b). colibre again shows very good convergence with resolution and is consistent with the observational data.